OpenStudy (anonymous):

if A+B =225, prove that : cot A /1+cot A , Cot b/1+cot B =1/2

OpenStudy (anonymous):

if: \[A+B = 225\] then: \[\frac{\cot(A)}{1+\cot(A)}\] \[\frac{\cot(B)}{1+\cot(B)}=\frac{1}{2}\] ? what about the first fraction? what does it equal?

OpenStudy (anonymous):

prove both are = to 1/2

OpenStudy (anonymous):

Where did you get this prblem from? its not necessarily true. If i take A = 180, and B = 45 (in degrees), i get: cot(180) = 0 cot(45) = 1 \[\frac{\cot(180)}{1+\cot(180)} = \frac{0}{1} = 0\] \[\frac{\cot(45)}{1+\cot(45)} = \frac{1}{2} \] Only one of them is equal to 1/2, not both.